# Reversible Deep Equilibrium Models

### v1

@@ Abstract
Deep Equilibrium Models (DEQs) are an interesting class of implicit model where the model output is implicitly defined as the fixed point of a learned function. These models have been shown to outperform explicit (fixed-depth) models in large-scale tasks by trading many deep layers for a single layer that is iterated many times. However, gradient calculation through DEQs is approximate. This often leads to unstable training dynamics and requires regularisation or many function evaluations to fix. Here, we introduce Reversible Deep Equilibrium Models (RevDEQs) that allow for exact gradient calculation, no regularisation and far fewer function evaluations than DEQs. We show that RevDEQs achieve state-of-the-art performance on language modelling and image classification tasks against comparable implicit and explicit models.
@@ 1 Introduction
Explicit models, where the forward model evaluation is explicitly defined and fixed, are the dominant modelling paradigm in machine learning. However, recent work has shown that it is possible to define the forward model implicitly, where a numerical algorithm is evaluated at runtime to determine the model output. This is instantiated by Neural Differential Equations chen2018neural ; gholaminejad2019anode ; kidger2020neural ; li2020scalable ; kidger2021neural ; issa2024non ; kidger2022neural , Optimisation Networks amos2017optnet ; gould2021deep and Deep Equilibrium Models (DEQs) bai2019deep ; el2021implicit .
Implicit models display interesting properties over their explicit counterparts: compute can be dynamically allocated per data example, memory usage is constant with respect to the effective model depth and oftentimes task performance is improved. Here, we focus on DEQs as the most general and performant class of implicit model with many applications: language bai2019deep , classification and segmentation bai2020multiscale , generative modelling bai2022deep ; lu2021implicit ; pokle2022deep , inverse problems gilton2021deep , implicit representations huang2021textrm and graph neural networks gu2020implicit ; liu2022mgnni .
However, a number of fundamental issues remain with the formulation of DEQs. Backpropagation is performed by computing a numerical solution to the implicit function theorem bai2019deep . The gradients obtained are therefore an approximation to the true gradient of the model being trained. This approximation causes training instability and oftentimes training diverges altogether bai2021stabilizing . To alleviate this, many function evaluations are required to compute an accurate gradient or strong regularisations are needed to ensure the gradient system is simpler to solve bai2021stabilizing ; winston2020monotone ; gabor2024positive .
@@ 2.1 Deep Equilibrium Models
A DEQ defines the model output, $\mathbf{y}$, as the fixed point of a parameterised function bai2019deep ,
@@
In practice, stronger fixed point solvers are used to accelerate convergence. Broyden’s method broyden1965class , a quasi-Newton root finding algorithm, was the original solver of choice bai2019deep . Recently, Anderson’s method anderson1965iterative has become popular due to the reduced computational cost and improved performance gilton2021deep .
@@ 2.2 Backpropagation
For DEQs, the gradient can be derived via the Implicit Function Theorem (IFT) bai2019deep ; winston2020monotone . Consider a scalar-valued loss function $L(z^{*}):\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}$ and associated gradient (adjoint) $\mathbf{a}_{z}=\partial L/\partial\mathbf{z}$. Then,
@@
Given the difficulty of training implicit models via the implicit function theorem, many alternative gradient approximations can be made: Neumann series lorraine2020optimizing ; geng2021training , phantom gradient geng2021training and Jacobian-free backpropagation fung2022jfb . Note that these methods introduce a further approximation to the exact gradient of the model but are cheaper to compute.
@@ 2.3 Algebraic Reversibility
We could simply store this graph in memory and backpropagate using standard automatic differentiation griewank2008evaluating . However, the memory cost of this graph scales with $N$. Instead, if our fixed point solver is algebraically reversible then the forward graph can be reconstructed exactly (up to floating point precision) during backpropagation. By reconstructing the forward graph exactly, gradient calculation becomes exact.
@@
Reversibility has been found to be a useful idea in many areas of machine learning: reversible networks apply this to residual networks gomez2017reversible ; chang2018reversible , ReFormer uses this construction in transformers kitaev2020reformer , momentum networks present an alternative reversible residual network sander2021momentum , coupling layers are used in generative modelling for exact likelihood calculation dinh2014nice and image editing wallace2023edict , reversible neural differential equation solvers improve backpropagation efficiency mccallum2024efficient ; kidger2021efficient ; zhuang2021mali .
@@ 3.4 Backpropagation
The reversible backpropagation algorithm proceeds by first reconstructing the state $\{\mathbf{y}_{n},\mathbf{z}_{n}\}$ from $\{\mathbf{y}_{n+1},\mathbf{z}_{n+1}\}$ using the backward step in equation (6). Then, the gradients are backpropagated by following the reverse-mode automatic differentiation rules griewank2008evaluating for the forward step in equation (5). See Algorithm 1.
We define a scalar-valued loss function $L(\mathbf{z}_{N}):\mathbb{R}^{d}\rightarrow\mathbb{R}$ on the terminal value of the solve $\mathbf{z}_{N}\approx\mathbf{z^{*}}$. The gradients (adjoints) are defined as $\bar{\mathbf{v}}=\partial L(\mathbf{z}_{N})/\partial\mathbf{v}$. The presence of an adjoint to the left of a Jacobian, $\bar{\mathbf{v}}\frac{\partial f_{\theta}}{\partial\mathbf{z}}$, indicates a vector-Jacobian product. This can be efficiently computed using automatic differentiation griewank2008evaluating ; griewank1992achieving .
@@ 4 Experiments
For language modelling, we use the Wikitext-103 dataset containing over 100 million tokens from the set of verified good and featured articles on Wikipedia merity2016pointer . We compare performance to previous DEQ models bai2019deep and an explicit transformer architecture via Transformer-XL dai2019transformerxl .
For image classification, we use the CIFAR-10 dataset krizhevsky2009learning . We introduce a novel multi-scale implicit architecture, inspired by the explicit ResNet model he2016deep , where the traditional deep unrolled convolutional layers at each image scale are replaced by a single RevDEQ block. This model is significantly simpler than previous multi-scale DEQ approaches bai2020multiscale ; winston2020monotone and achieves higher classification accuracy. We compare performance to previous single-scale and multi-scale implicit models, and explicit ResNet architectures.
@@ 4.1 Language Modelling on Wikitext-103
The original DEQ model was applied to model sequential data, such as language bai2019deep . This was motivated by the observation that the hidden layers of explicit deep architectures naturally learn contractive functions and push the hidden state towards a fixed point.
We evaluate the performance of the RevDEQ model on the Wikitext-103 dataset merity2016pointer . The model is parameterised with a decoder-only transformer architecture vaswani2017attention , but with the many explicit layers replaced by a single layer, $f_{\theta}$, of which we find the fixed point.
@@
First the input sequence is added to the hidden state $\mathbf{z}_{1:T}$, then we apply a self-attention operation with $W_{QKV}\in\mathbb{R}^{3d\times d}$ vaswani2017attention followed by layer normalisation (LN) ba2016layer , 2-layer MLP $\phi$, and a final LN. The self-attention operation is multi-headed, with $12$ heads used for the experiments below.
The fixed point of $f_{\theta}$ is calculated using the reversible fixed point solver in equation (5). Gradient calculation is performed via the reversible backpropagation algorithm, given by Algorithm 1. We set the fixed point tolerance to be $10^{-3}$ and choose a maximum of 4 solver steps. See Appendix B for full details.
Two models sizes are tested: small (110M parameters) and medium (169M parameters). We compare the performance to comparable implicit and explicit models in Table 1.
@@
Transformer-XL (4 layers) dai2019transformerxl 139M - 35.8
DEQ bai2019deep 138M 30 32.4
@@
Transformer-XL (16 layers) dai2019transformerxl 165M - 24.3
DEQ-TrellisNet bai2019deep 180M 30 29.0
DEQ-Transformer bai2019deep 172M 30 24.2
RevDEQ 169M 8 20.7
Table 1: Test perplexity on the Wikitext-103 dataset. RevDEQ models outperform previous DEQ approaches while requiring far fewer function evaluations. The RevDEQ model also outperforms explicit transformer architectures of comparable size.
@@
We hypothesise that the improved performance is largely a function of exact gradient calculation. To test this, we vary the number of function evaluations in the 169M parameter RevDEQ model and calculate test perplexity. This is shown in Table 2.
It can be seen that the RevDEQ model is competitive with previous DEQ models when using only 2 function evaluations. Additionally, the perplexity quickly decreases as the number of function evaluations increase and plateaus after approximately 10 evaluations. This scaling is much faster than previous implicit models, where typically 10s or 100s of evaluations are required to reach optimality.
@@
This result can be understood by considering the following argument. The explicit transformer architecture distributes the learnable parameters over $N$ layers. But it is known that an $N$ layer deep network can be represented by a single weight-tied layer that is iterated for (at most) $N$ steps, where the single hidden state is formed by the concatenation of the hidden states over the $N$ layers (see Theorem 3 of bai2019deep ).
@@
DEQ-Transformer bai2019deep 30 24.2
DEQ-Transformer (JR) bai2021stabilizing 14 24.5
RevDEQ 2 24.6
4 22.0
6 21.3
8 20.7
10 20.8
Table 2: Test perplexity on the Wikitext-103 dataset in terms of the number of function evaluations. RevDEQ models are more compute efficient in comparison to previous DEQs. JR stands for Jacobian Regularisation.
@@ 4.2 Image Classification on CIFAR-10
Previous multi-scale implicit architectures have focused on concatenating (width-wise) all scales into one feature vector where each scale is driven to equilibrium in parallel bai2020multiscale ; winston2020monotone . Here, we instead choose to model each image scale with an implicit architecture and parameterise explicit downsampling layers between scales. This is a ResNet model he2016deep , but with the many deep unrolled convolutional blocks at each image scale replaced by a single convolutional block of which we find the fixed point.
This parameterisation is much simpler than previous multi-scale architectures bai2020multiscale ; winston2020monotone ; additionally, we find the model to be more parameter efficient and performant, even in comparison to explicit ResNet models he2016deep .
@@
This is iterated four times, defining four feature scales as is common in ResNet models he2016deep . At each scale, $i$, the number of channels $C_{i}$ is selected along with the fixed point tolerance $\epsilon_{i}$ and maximum steps $N_{i}$. The number of steps $N_{i}$ defines the effective depth of the model at each scale.
The downsampling layers follow the design of he2016identity , with the image dimension $(H_{i},W_{i})$ halved by a strided convolution operation. The residual connection follows a 1x1 convolution to align the channel dimensions $C_{i}\rightarrow C_{i+1}$. Finally, we apply the linear classification layers. See Appendix B for full details.
@@ Results
DEQ bai2019deep 170K 15 $82.2\pm 0.3$
monDEQ winston2020monotone 854K 20 $82.0\pm 0.3$
@@
MDEQ bai2020multiscale 170K 15 $87.1\pm 0.4$
monDEQ winston2020monotone 1M 20 $89.0\pm 0.3$
pcDEQ gabor2024positive 661K 14 $89.2\pm 0.0$
@@
ResNet-18 he2016deep 10M - $92.9\pm 0.2$
ResNet-101 he2016deep 40M - $93.8\pm 0.3$
MDEQ bai2020multiscale 10M 15 $93.8\pm 0.2$
MDEQ (JR) bai2021stabilizing 10M 6 $93.1\pm 0.3$
@@ 5.2 Limitations
The normalisations used inside the equilibrium function $f_{\theta}$ are stateless layer-wise. This is due to the number of layers in the model being defined implicitly, making it challenging to collect population statistics for each ‘layer’. In the implicit ResNet model, to obtain stateful normalisation over the hidden state we use batch normalisation ioffe2015batch in the downsampling blocks prior to a RevDEQ block.
@@ 5.3 Future work
There are many applications of implicit architectures in a diverse set of tasks, such as graph neural networks gu2020implicit , flow lu2021implicit ; bai2022deep and diffusion models pokle2022deep , implicit neural representations huang2021textrm and inverse problems gilton2021deep . The RevDEQ model is yet to be applied to these problems!
@@
The arithmetic intensity of modern GPUs is typically $\sim 10^{2}$, meaning that if less than $10^{2}$ FLOPs are performed per byte of data then algorithm speed is limited by memory operations scaling-book . In standard backpropagation, many read/write operations are required to store and load the forward pass of the network griewank2008evaluating ; dao2022flashattention . For implicit models, this is not required - completely removing these memory operations. By careful algorithm design, this could be exploited to improve the GPU efficiency of implicit models.

### v2

@@ Abstract
Deep Equilibrium Models (DEQs) are an interesting class of implicit model where the model output is implicitly defined as the fixed point of a learned function. These models have been shown to outperform explicit (fixed-depth) models in large-scale tasks by trading many deep layers for a single layer that is iterated many times. However, gradient calculation through DEQs is approximate. This often leads to unstable training dynamics and requires regularisation or many function evaluations to fix. Here, we introduce Reversible Deep Equilibrium Models (RevDEQs) that allow for exact gradient calculation, no regularisation and far fewer function evaluations than DEQs. We show that RevDEQs significantly improve performance on language modelling and image classification tasks against comparable implicit and explicit models.
@@ 1 Introduction
Explicit models, where the forward model evaluation is explicitly defined and fixed, are the dominant modelling paradigm in machine learning. However, recent work has shown that it is possible to define the forward model implicitly, where a numerical algorithm is evaluated at runtime to determine the model output. This is instantiated by Neural Differential Equations [chen2018neural, gholaminejad2019anode, kidger2020neural, li2020scalable, kidger2021neural, issa2024non, kidger2022neural], Optimisation Networks [amos2017optnet, gould2021deep] and Deep Equilibrium Models (DEQs) [bai2019deep, el2021implicit].
Implicit models display interesting properties over their explicit counterparts: compute can be dynamically allocated per data example, memory usage is constant with respect to the effective model depth and oftentimes task performance is improved. Here, we focus on DEQs as the most general and performant class of implicit model with many applications: language [bai2019deep], classification and segmentation [bai2020multiscale], generative modelling [bai2022deep, lu2021implicit, pokle2022deep], inverse problems [gilton2021deep], implicit representations [huang2021textrm] and graph neural networks [gu2020implicit, liu2022mgnni].
However, a number of fundamental issues remain with the formulation of DEQs. Backpropagation is performed by computing a numerical solution to the implicit function theorem [bai2019deep]. The gradients obtained are therefore an approximation to the true gradient of the model being trained. This approximation causes training instability and oftentimes training diverges altogether [bai2021stabilizing]. To alleviate this, many function evaluations are required to compute an accurate gradient or strong regularisations are needed to ensure the gradient system is simpler to solve [bai2021stabilizing, winston2020monotone, gabor2024positive].
@@ 2.1 Deep Equilibrium Models
A DEQ defines the model output, $\mathbf{y}$, as the fixed point of a parameterised function [bai2019deep],
@@
In practice, stronger fixed point solvers are used to accelerate convergence. Broyden’s method [broyden1965class], a quasi-Newton root finding algorithm, was the original solver of choice [bai2019deep]. Recently, Anderson’s method [anderson1965iterative] has become popular due to the reduced computational cost and improved performance [gilton2021deep].
@@ 2.2 Backpropagation
For DEQs, the gradient can be derived via the Implicit Function Theorem (IFT) [bai2019deep, winston2020monotone]. Consider a scalar-valued loss function $L(z^{*}):\mathbb{R}^{d_{z}}\rightarrow\mathbb{R}$ and associated gradient (adjoint) $\mathbf{a}_{z}=\partial L/\partial\mathbf{z}$. Then,
@@
Given the difficulty of training implicit models via the implicit function theorem, many alternative gradient approximations can be made: Neumann series [lorraine2020optimizing, geng2021training], phantom gradient [geng2021training] and Jacobian-free backpropagation [fung2022jfb]. Note that these methods introduce a further approximation to the exact gradient of the model but are cheaper to compute.
@@ 2.3 Algebraic Reversibility
We could simply store this graph in memory and backpropagate using standard automatic differentiation [griewank2008evaluating]. However, the memory cost of this graph scales with $N$. Instead, if our fixed point solver is algebraically reversible then the forward graph can be reconstructed exactly (up to floating point precision) during backpropagation. By reconstructing the forward graph exactly, gradient calculation becomes exact.
@@
Reversibility has been found to be a useful idea in many areas of machine learning: reversible networks apply this to residual networks [gomez2017reversible, chang2018reversible], ReFormer uses this construction in transformers [kitaev2020reformer], momentum networks present an alternative reversible residual network [sander2021momentum], coupling layers are used in generative modelling for exact likelihood calculation [dinh2014nice] and image editing [wallace2023edict], reversible neural differential equation solvers improve backpropagation efficiency [mccallum2024efficient, kidger2021efficient, zhuang2021mali].
@@ 3.4 Backpropagation
The reversible backpropagation algorithm proceeds by first reconstructing the state $\{\mathbf{y}_{n},\mathbf{z}_{n}\}$ from $\{\mathbf{y}_{n+1},\mathbf{z}_{n+1}\}$ using the backward step in equation (6). Then, the gradients are backpropagated by following the reverse-mode automatic differentiation rules [griewank2008evaluating] for the forward step in equation (5). See Algorithm 1.
We define a scalar-valued loss function $L(\mathbf{z}_{N}):\mathbb{R}^{d}\rightarrow\mathbb{R}$ on the terminal value of the solve $\mathbf{z}_{N}\approx\mathbf{z^{*}}$. The gradients (adjoints) are defined as $\bar{\mathbf{v}}=\partial L(\mathbf{z}_{N})/\partial\mathbf{v}$. The presence of an adjoint to the left of a Jacobian, $\bar{\mathbf{v}}\frac{\partial f_{\theta}}{\partial\mathbf{z}}$, indicates a vector-Jacobian product. This can be efficiently computed using automatic differentiation [griewank2008evaluating, griewank1992achieving].
@@ 4 Experiments
For language modelling, we use the Wikitext-103 dataset containing over 100 million tokens from the set of verified good and featured articles on Wikipedia [merity2016pointer]. We compare performance to previous DEQ models [bai2019deep, bai2021stabilizing] and an explicit transformer architecture via Transformer-XL [dai2019transformerxl].
For image classification, we use the CIFAR-10 dataset [krizhevsky2009learning]. We introduce a novel multi-scale implicit architecture, inspired by the explicit ResNet model [he2016deep], where the traditional deep unrolled convolutional layers at each image scale are replaced by a single RevDEQ block. This model is significantly simpler than previous multi-scale DEQ approaches [bai2020multiscale, winston2020monotone] and achieves higher classification accuracy. We compare performance to previous single-scale and multi-scale implicit models, and explicit ResNet architectures.
@@ 4.1 Language Modelling on Wikitext-103
The original DEQ model was applied to model sequential data, such as language [bai2019deep]. This was motivated by the observation that the hidden layers of explicit deep architectures naturally learn contractive functions and push the hidden state towards a fixed point.
We evaluate the performance of the RevDEQ model on the Wikitext-103 dataset [merity2016pointer]. The model is parameterised with a decoder-only transformer architecture [vaswani2017attention], but with the many explicit layers replaced by a single layer, $f_{\theta}$, of which we find the fixed point.
@@
First the input sequence is added to the hidden state $\mathbf{z}_{1:T}$, then we apply a self-attention operation with $W_{QKV}\in\mathbb{R}^{3d\times d}$ [vaswani2017attention] followed by layer normalisation (LN) [ba2016layer], 2-layer MLP $\phi$, and a final LN. The self-attention operation is multi-headed, with $8$ heads used for the experiments below.
The fixed point of $f_{\theta}$ is calculated using the reversible fixed point solver in equation (5). Gradient calculation is performed via the reversible backpropagation algorithm, given by Algorithm 1. We set the fixed point tolerance to be $10^{-3}$ and choose a maximum number of solver steps. See Appendix B for full details.
We compare RevDEQ performance to comparable implicit and explicit models in Table 1. This includes the Transformer-XL explicit architecture [dai2019transformerxl], the original DEQ-Transformer [bai2019deep] and the DEQ-Transformer with Jacobian regularisation (JR) [bai2021stabilizing].
@@
Transformer-XL (18 layers) [dai2019transformerxl] 110M - 24.1
DEQ-Transformer [bai2019deep] 98M 30 24.0
DEQ-Transformer + JR [bai2021stabilizing] 98M 14 24.5
DEQ-Transformer + JR [bai2021stabilizing] 98M 12 24.9
@@
RevDEQ 94M 12 23.2
Table 1: Test perplexity on the Wikitext-103 dataset. RevDEQ models outperform previous DEQ approaches while requiring fewer function evaluations. The RevDEQ model also outperforms explicit transformer architectures of comparable size.
@@
We hypothesise that the improved performance is largely a function of exact gradient calculation. To test this, we vary the number of function evaluations in the RevDEQ model and calculate test perplexity. This is shown in Table 2.
It can be seen that the RevDEQ model is competitive with previous DEQ models when using only 6 function evaluations. Additionally, the perplexity quickly decreases as the number of function evaluations increase and plateaus after approximately 12 evaluations. This scaling is much faster than previous implicit models, where typically 10s or 100s of evaluations are required to reach optimality.
@@
This result can be understood by considering the following argument. The explicit transformer architecture distributes the learnable parameters over $N$ layers. But it is known that an $N$ layer deep network can be represented by a single weight-tied layer that is iterated for (at most) $N$ steps, where the single hidden state is formed by the concatenation of the hidden states over the $N$ layers (see Theorem 3 of [bai2019deep]).
@@
DEQ-Transformer [bai2019deep] 30 24.0
DEQ-Transformer (JR) [bai2021stabilizing] 12 24.9
RevDEQ 2 27.3
4 24.6
6 23.7
8 23.4
10 23.3
12 23.2
Table 2: Test perplexity on the Wikitext-103 dataset in terms of the number of function evaluations. RevDEQ models are more compute efficient in comparison to previous DEQs.
@@ 4.2 Image Classification on CIFAR-10
Previous multi-scale implicit architectures have focused on concatenating (width-wise) all scales into one feature vector where each scale is driven to equilibrium in parallel [bai2020multiscale, winston2020monotone]. Here, we instead choose to model each image scale with an implicit architecture and parameterise explicit downsampling layers between scales. This is a ResNet model [he2016deep], but with the many deep unrolled convolutional blocks at each image scale replaced by a single convolutional block of which we find the fixed point.
This parameterisation is much simpler than previous multi-scale architectures [bai2020multiscale, winston2020monotone]; additionally, we find the model to be more parameter efficient and performant, even in comparison to explicit ResNet models [he2016deep].
@@
This is iterated four times, defining four feature scales as is common in ResNet models [he2016deep]. At each scale, $i$, the number of channels $C_{i}$ is selected along with the fixed point tolerance $\epsilon_{i}$ and maximum steps $N_{i}$. The number of steps $N_{i}$ defines the effective depth of the model at each scale.
The downsampling layers follow the design of [he2016identity], with the image dimension $(H_{i},W_{i})$ halved by a strided convolution operation. The residual connection follows a 1x1 convolution to align the channel dimensions $C_{i}\rightarrow C_{i+1}$. Finally, we apply the linear classification layers. See Appendix B for full details.
@@ Results
DEQ [bai2019deep] 170K 15 $82.2\pm 0.3$
monDEQ [winston2020monotone] 854K 20 $82.0\pm 0.3$
@@
MDEQ [bai2020multiscale] 170K 15 $87.1\pm 0.4$
monDEQ [winston2020monotone] 1M 20 $89.0\pm 0.3$
pcDEQ [gabor2024positive] 661K 14 $89.2\pm 0.0$
@@
ResNet-18 [he2016deep] 10M - $92.9\pm 0.2$
ResNet-101 [he2016deep] 40M - $93.8\pm 0.3$
MDEQ [bai2020multiscale] 10M 15 $93.8\pm 0.2$
MDEQ (JR) [bai2021stabilizing] 10M 6 $93.1\pm 0.3$
@@ 5.2 Limitations
The normalisations used inside the equilibrium function $f_{\theta}$ are stateless layer-wise. This is due to the number of layers in the model being defined implicitly, making it challenging to collect population statistics for each ‘layer’. In the implicit ResNet model, to obtain stateful normalisation over the hidden state we use batch normalisation [ioffe2015batch] in the downsampling blocks prior to a RevDEQ block.
@@ 5.3 Future work
There are many applications of implicit architectures in a diverse set of tasks, such as graph neural networks [gu2020implicit], flow [lu2021implicit, bai2022deep] and diffusion models [pokle2022deep], implicit neural representations [huang2021textrm] and inverse problems [gilton2021deep]. The RevDEQ model is yet to be applied to these problems!
@@
The arithmetic intensity of modern GPUs is typically $\sim 10^{2}$, meaning that if less than $10^{2}$ FLOPs are performed per byte of data then algorithm speed is limited by memory operations [scaling-book]. In standard backpropagation, many read/write operations are required to store and load the forward pass of the network [griewank2008evaluating, dao2022flashattention]. For implicit models, this is not required - completely removing these memory operations. By careful algorithm design, this could be exploited to improve the GPU efficiency of implicit models.